# Space-time first-order correlations of an open Bose Hubbard model with incoherent pump and loss

### v1

@@ Abstract
We investigate the correlation properties in the steady state of driven-dissipative interacting bosonic systems in the quantum regime, as for example non-linear photonic cavities. Specifically, we consider the Bose-Hubbard model on a periodic chain and with spatially homogeneous one-body loss and pump within the Markovian approximation. The steady state corresponds to an infinite temperature state at finite chemical potential with diagonal spatial correlations. Nonetheless, we observe a nontrivial behaviour of the space-time two-point correlation function in the steady state, obtained by exact diagonalisation. In particular, we find that the decay width of the propagator is not only renormalised at increasing interactions, as it is the case of a single non-linear resonator, but also at increasing hopping strength. We then compute the full spectral function, finding that it contains both a dispersive free-particle like dispersion at low energy and a doublon branch at energy corresponding to the on-site interactions. We compare with the corresponding calculation for the ground state of a closed quantum system and show that the driven-dissipative nature – determining both the steady state and the dynamical evolution – changes the low-lying part of the spectrum, where noticeably, the dispersion is quadratic instead of linear at small wavevectors.
@@ 1 Introduction
While all these studies concern closed quantum systems, in several experimental situations quantum systems are subjected to external pump and/or losses, or put in contact with some type of bath. Examples of bosonic open systems are Josephson junction arrays [25], superconducting microwave circuits [26], polariton chains [27], atoms or ions in optical cavities  [28, 29], optomechanical systems [30], and lossy quantum gases [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. In contrast to the equilibrium case, much less is known for interacting open quantum systems, and several questions arise, ranging from the properties of the steady state and its phase diagram to its coherence, its excitations and its dynamical properties. It is also very interesting to investigate which properties known for unitary systems are still present in open quantum systems, and under which conditions or parameters.
@@
A way to model an open system is to describe it as a subsystem embedded in a larger system acting as a bath and with which it interacts. If one requires the time evolution of the subsystem to be completely-positive and trace-preserving (CPTP map), this evolution has to be of the Lindblad form [42], and described by a Markovian quantum master equation. This evolution follows the composition law for universal dynamical maps which is the quantum analogue to the classical differential Chapman-Kolmogorov equation [43]. Despite its simplicity, the Lindblad master equation suffices to display rich physical behaviour, as it has been demonstrated for the case of the steady state of interacting driven-dissipative bosons [44, 45]. Quadratic Lindbladians, meaning Lindbladians with quadratic Hamiltonian and linear jump operators in the bosonic (fermionic) creation and annihilation operators, are known to be exactly solvable by the method of third quantisation, presented by Prosen [46] and Prosen and Seligman [47]. Within the Keldysh formalism these systems correspond to a quadratic action. Also instances of quasi-free Lindbladians for fermions and bosons with quadratic hermitean jump operators have been solved [48, 49]. Recently, the connection between both formalisms and the phase-space formulation has been pioneered [50]. Non-quadratic Lindbladians require in most cases a numerical solution. We refer to [51] for a comparative analysis of the state-of-art numerical methods, specifically tailored for non-equal time correlation functions. This method was used to obtain the second-order correlation function $g^{(2)}$ [52].
@@
An especially important quantity to characterise a quantum system is its two-point correlation function, corresponding to first-order correlations $g^{(1)}$, see e.g. [53]. At equal times, it contains information about the spatial coherence and the presence of (quasi)off-diagonal long-range order, i.e. the existence of Bose-Einstein condensation according to the Penrose Onsager criterion [54]. Its Fourier transform yields the momentum distribution function, giving information on the velocity of the quantum particles in a given quantum state. The full space-time correlation function describes the evolution of the quantum system when removing or adding a particle to it. In space-time, it allows one to study the spread of correlations, and to test the Lieb-Robinson bound [55], and its generalisation to open quantum systems [56]. Its Fourier transform in frequency-wavevector space provides the spectral function, whose poles give the dispersion of single-particle excitations. The spectral function is also needed to describe the transport properties among two reservoirs across a quantum channel within linear response [57].
@@
In this work, we study the two-point correlation function of an interacting bosonic open quantum system described by the Bose-Hubbard model subjected to incoherent drive and losses. At non-zero interactions, similarly to its closed-system counterpart, this model is not integrable. We choose the case of spatially uniform pump and losses, where the non-equilibrium steady-state (NESS) density matrix is known exactly [58], and it is closely related to an infinite-temperature state at fixed chemical potential, independently of the interaction strength. Despite the steady state taking a simple form, we find that the two-point correlation function shows a non-trivial dependence on the interaction strength, displaying in particular typical temporal oscillations with frequency related to the interaction strength as well as an interaction-dependent exponential decay. We compare these results to the exactly known non-interacting limit as well as to the strongly interacting limit where the dynamics of the different sites decouples and can be described using the exact results for the dissipative Kerr resonator [59]. We then proceed in obtaining the full spectral function, where in particular we find a doublon-like branch at excitation energy $U$ as well as a cosine-like excitation branch in the single-particle-hole sector. The features of the spectral function are a clear signature of the non-equilibrium nature of the quantum system, further supported by a Keldysh action analysis, and by the comparison with the closed-quantum system spectral function [60].
@@
The paper is organised as follows. We start by presenting the model and the characterisation of the non-equilibrium steady state in Sec. 2, which includes the Lindblad equation in Sec. 2.1 and the Keldysh action in Sec. 2.3. We first provide our predictions for the particle density as a function of time in Sec. 3. We then proceed to present our results for the space-time correlations in Sec. 4, where we display the time-resolved correlations and analyse their decay. The spectral function is shown in Sec. 5. Finally, Sec. 6 presents our concluding remarks.
@@ 2.1 Lindblad equation
$H=\sum_{i=1}^{L}\Big{[}-J\big{(}b_{i}^{\dagger}b_{i+1}+\text{h.c.}\big{)}+\frac{U}{2}b_{i}^{\dagger}b_{i}^{\dagger}b_{i}b_{i}\Big{]}\;,$ (1)
@@
$\partial_{t}\rho=\mathcal{L}[\rho]=-\text{i}[H,\rho]+\gamma_{l}\sum_{i=1}^{L}\big{(}b_{i}\rho b_{i}^{\dagger}-\frac{1}{2}\{b_{i}^{\dagger}b_{i},\rho\}\big{)}+\gamma_{p}\sum_{i=1}^{L}\big{(}b_{i}^{\dagger}\rho b_{i}-\frac{1}{2}\{b_{i}b_{i}^{\dagger},\rho\}\big{)}\;.$ (2)
@@ 2.2 Correlation functions
where delimited-⟨$\langle$..$\rangle$ … ⟩ indicates the trace weighted with the respective density matrix, $\theta(t)$ is the Heavyside step function, and $b_{j}(t)$ ($b^{\dagger}_{j}(t)$) is the bosonic annihilation (creation) field operator in the Heisenberg picture, time-evolved according to Lindbladian dynamics [^1] In the case of open quantum systems, a way of defining the time evolution of an operator $A$ is given by the adjoint superoperator $\bar{\mathcal{L}}$, according to $A(t)=e^{\bar{\mathcal{L}}t}[A]$ where ¯$\bar{\mathcal{L}}[A]=+\text{i}[H,A]+\gamma_{l}\sum_{i=1}^{L}\big{(}b_{i}^{\dagger}Ab_{i}-\frac{1}{2}\{b_{i}^{\dagger}b_{i},A\}\big{)}+\gamma_{p}\sum_{i=1}^{L}\big{(}b_{i}Ab_{i}^{\dagger}-\frac{1}{2}\{b_{i}b_{i}^{\dagger},A\}\big{)}$. Notice that ${\rm Tr}\big{\{}A(t)\rho(0)\big{\}}={\rm Tr}\big{\{}A\rho(t)\big{\}}$. (in the open case), or Hamiltonian one (in the closed case).
@@
For a closed system the corresponding correlation function evaluated on the ground state $|\Psi_{0}\rangle$ reads
$\Big{\langle}b_{j}^{\dagger}(t)b_{0}\Big{\rangle}_{\text{GS}}\equiv\langle\Psi_{0}|\text{e}^{\text{i}Ht}\,b_{j}^{\dagger}\text{e}^{-\text{i}Ht}\,b_{0}|\Psi_{0}\rangle.$ (6)
@@
for a homogeneous system, and with the Fourier conventions
[…]
$f_{j}(t)$ $=\dfrac{1}{L}\int_{-\infty}^{\infty}\dfrac{dp}{2\pi}\int_{-\infty}^{\infty}\dfrac{d\omega}{2\pi}\,f(p,\omega)\,e^{-ipj}e^{i\omega t}\,.$ (8)
2.3 Keldysh field theory
It is useful to formulate the model in terms of the Keldysh formalism, following the derivation of [45]. Starting from the Lindblad equation (2), and expanding the system density matrix onto two coherent-state bases for each space-time point identified by the fields $\Psi_{j}^{+}(t)$ ($\Psi_{j}^{-}(t)$) for upper (lower) time contours, one obtains the time evolution of the density matrix in terms of a path integral. Following the usual notations, the Keldysh action in the rotated basis $({\Psi_{c}}_{j}(t),{\Psi_{q}}_{j}(t))=\frac{1}{\sqrt{2}}(\Psi_{j}^{+}(t)+\Psi_{j}^{-}(t),\Psi_{j}^{+}(t)-\Psi_{j}^{-}(t))$ reads
$S=\int\text{d}t\sum_{j}\Bigg{[}$ ${\Psi_{c}}_{j}^{*}\,\text{i}\left(\partial_{t}-\frac{\gamma_{l}-\gamma_{p}}{2}\right){\Psi_{q}}_{j}+{\Psi_{q}}_{j}^{*}\,\text{i}\left(\partial_{t}+\frac{\gamma_{l}-\gamma_{p}}{2}\right){\Psi_{c}}_{j}+\text{i}(\gamma_{l}+\gamma_{p})|{\Psi_{q}}_{j}|^{2}$
$+J\Big{(}{\Psi_{c}}_{j+1}^{*}{\Psi_{q}}_{j}+{\Psi_{q}}_{j+1}^{*}{\Psi_{c}}_{j}+{\Psi_{c}}_{j}^{*}{\Psi_{q}}_{j+1}+{\Psi_{q}}_{j}^{*}{\Psi_{c}}_{j+1}\Big{)}$
$-\frac{U}{2}\Big{(}({\Psi_{c}}_{j}^{2}+{\Psi_{q}}_{j}^{2}){\Psi_{c}}_{j}^{*}{\Psi_{q}}_{j}^{*}+({{\Psi_{c}}_{j}^{*}}^{2}+{{\Psi_{q}}_{j}^{*}}^{2}){\Psi_{c}}_{j}{\Psi_{q}}_{j}\Big{)}\Bigg{]}\;.$ (9)
In the same notation, the retarded Green’s function (3) reads
$G^{R}_{j\ell}(t,t^{\prime})=-\text{i}\Big{\langle}\Psi_{cj}(t)\Psi_{q\ell}^{*}(t^{\prime})\Big{\rangle}_{c}\,.$ (10)
The saddle-point approximation in the path integral with the action (2.3) allows one to study the existence and stability of a condensate. For a static homogeneous field solution, if $\gamma_{p}<\gamma_{l}$, the imaginary part of the Keldysh potential has a global minimum for $\Psi_{cj}=\Psi_{qj}=0$, which implies that the system is in the symmetric phase without breaking of the U(1) symmetry. Hence, there is no finite order parameter. As we shall see below, this is in agreement with the exact NESS solution, which rather corresponds to an infinite-temperature state for the bosons. Notice that this is at variance with the case where both one- and two-body losses are present, where a U(1) symmetry-broken phase emerges for $\gamma_{p}>\gamma_{l}$ and a (quasi-)condensate forms at weak interactions [62].
Expanding the action around this solution to second order in the fluctuations, we obtain the inverse of the retarded sector of the propagator as
$[G_{0}^{R}]^{-1}_{ml}(t^{\prime},t)$ $\equiv\begin{pmatrix}\frac{\delta^{2}S}{\delta{\Psi^{*}_{qm}}(t^{\prime}){\delta\Psi_{cl}}(t)}&&\frac{\delta^{2}S}{\delta{\Psi^{*}_{qm}}(t^{\prime}){\delta\Psi^{*}_{cl}}(t)}\\ \frac{\delta^{2}S}{\delta{\Psi_{qm}}(t^{\prime}){\delta\Psi_{cl}}(t)}&&\frac{\delta^{2}S}{\delta{\Psi_{qm}}(t^{\prime}){\delta\Psi^{*}_{cl}}(t)}\end{pmatrix}\Bigg{|}_{\Psi_{c}=\Psi_{q}=0}$
$=\begin{pmatrix}(\text{i}\partial_{t}+\text{i}\kappa_{0})\delta_{m,l}+J(\delta_{m,l+1}+\delta_{m,l-1})&&0\\ 0&&-(\text{i}\partial_{t}+\text{i}\kappa_{0})\delta_{m,l}+J(\delta_{m,l+1}+\delta_{m,l-1})\end{pmatrix}$
×$\times\delta(t-t^{\prime})\;,$ (11)
where we defined
$\kappa_{0}=(\gamma_{l}-\gamma_{p})/2,$ (12)
and where the 0 index indicates that this is the ‘bare’ propagator, obtained within the quadratic (non-interacting) approximation. In order to calculate the excitation branches $\Omega(k)$ in this approximation, we search for the solutions implicitly given by [45, 63]
$[G_{0}^{R}]^{-1}(k,\Omega(k))\begin{pmatrix}\Psi_{c}\\ \Psi^{*}_{c}\end{pmatrix}=0\;.$ (13)
By Fourier transforming Eq. (11), we find
$\Omega_{\pm}(k)=-\text{i}\kappa_{0}\mp 2J\cos(k)\;.$ (14)
This approach allows us to identify $\kappa_{0}$ with the decay rate at weak interactions corresponding to the broadening of the spectral line, as well as the excitation dispersion $\Omega_{+}$, corresponding to the free-particle dispersion in a lattice. Inverting the matrix (11) in Fourier space and selecting the relevant element (10), we obtain the retarded Green’s function
$G_{0}^{R}(k,\omega)=\frac{1}{\omega+2J\cos k+\text{i}\kappa_{0}}\;,$ (15)
which has poles for $\omega=\Omega_{+}(k)$. Notice that the second branch $\Omega_{-}(k)$ has zero spectral weight in this approximation. Hence, in the open system under investigation, there is no ghost branch predicted at quadratic order in the fluctuations, due to the absence of a quasi Bose-Einstein condensate. To include the effect of interactions, we turn to numerical simulations in Secs. 4 and 5 below. Prior to this, we present the steady-state density matrix.
@@ 3 Exact properties of the non-equilibrium steady state
3.1 Steady state density matrix
@@
We remark that, in this idealised model, there is no high-energy cutoff on the pump intensity. We refer to Ref. [58] for a model with frequency dependent effective jump operators. The infinite dimensional Hilbert space of the model as written above renders the Hamiltonian unbounded. We point out that in the numerical calculations, we have to introduce a cutoff for the local occupation, which gives an upper bound for the Hamiltonian and ensures that the time evolution is a CPT map as well as the existence of a NESS [43]. We have generalised the solution (16) for $\rho_{\text{NESS}}$ to the case of driven-dissipative free fermions and hard-core bosons on the lattice. The results are given in Appendix C.

### v2

@@ Abstract
We investigate the correlation properties in the steady state of driven-dissipative interacting bosonic systems in the quantum regime, as for example non-linear photonic cavities. Specifically, we consider the Bose-Hubbard model on a periodic chain and with spatially homogeneous one-body loss and pump within the Markovian approximation. The steady state is non-thermal and is formally equivalent to an infinite-temperature state with finite chemical potential set by the dissipative parameters. While there is no effect of interactions on the steady state, we observe a nontrivial behaviour of the space-time two-point correlation function, obtained by exact diagonalisation. In particular, we find that the decay width of the propagator is not only renormalised at increasing interactions, as it is the case of a single non-linear resonator, but also at increasing hopping strength. Furthermore, we numerically predict at large interactions a plateau value of the decay rate which goes beyond perturbative results in the interaction strength. We then compute the full spectral function, finding that it contains both a dispersive free-particle like dispersion at low energy and a doublon branch at energy corresponding to the on-site interactions. We compare with the corresponding calculation for the ground state of a closed quantum system and show that the driven-dissipative nature – determining both the steady state and the dynamical evolution – changes the low-lying part of the spectrum, where noticeably, the dispersion is quadratic instead of linear at small wavevectors. Finally, we compare to a high temperature grand-canonical equilibrium state and show the difference with respect to the open system stemming from the additional degree of freedom of the dissipation that allows one to vary the width of the dispersion lines.
@@ 1 Introduction
While all these studies concern closed quantum systems, in several experimental situations quantum systems are subjected to external pump and/or losses, or put in contact with some type of bath. Examples of bosonic open systems are Josephson junction arrays [25], superconducting microwave circuits [26, 27, 28], polariton chains [29], excitonic quantum dots coupled to cavity arrays [30, 31, 32], atoms or ions in optical cavities [33, 34], optomechanical systems [35], driven-dissipative [36] and lossy quantum gases [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. Ultracold bosonic atoms in transport geometries also realise non-equilibrium states (for reviews see e.g. [49, 50]). In contrast to the equilibrium case, much less is known for interacting open quantum systems, and several questions arise, ranging from the properties of the steady state and its phase diagram to its coherence, its excitations and its dynamical properties. It is also very interesting to investigate which properties known for unitary systems are still present in open quantum systems, and under which conditions or parameters.
@@
A way to model an open system is to describe it as a subsystem embedded in a larger system acting as a bath and with which it interacts. If one requires the time evolution of the subsystem to be completely-positive and trace-preserving (CPTP map), this evolution has to be of the Lindblad form [51], and described by a Markovian quantum master equation. This evolution follows the composition law for universal dynamical maps which is the quantum analogue to the classical differential Chapman-Kolmogorov equation [52]. Despite its simplicity, the Lindblad master equation suffices to display rich physical behaviour, as it has been demonstrated for the case of the steady state of interacting driven-dissipative bosons [53, 54]. Quadratic Lindbladians, meaning Lindbladians with quadratic Hamiltonian and linear jump operators in the bosonic (fermionic) creation and annihilation operators, are known to be exactly solvable by the method of third quantisation, presented by Prosen [55] and Prosen and Seligman [56]. Within the Keldysh formalism, these systems correspond to a quadratic action. Moreover, instances of quasi-free Lindbladians for fermions and bosons with quadratic hermitian jump operators have been solved [57, 58]. Recently, the connection between both formalisms and the phase-space formulation has been pioneered [59]. Non-quadratic Lindbladians require in most cases a numerical solution (see e.g. [60]). We refer to [61] for a comparative analysis of the state-of-art numerical methods, specifically tailored for non-equal time correlation functions. This method was used to obtain the second-order correlation function $g^{(2)}$ [62].
@@
A particularly important quantity to characterise a quantum system is its two-point correlation function, corresponding to first-order correlations $g^{(1)}$, see e.g. [63]. At equal times, it contains information about the spatial coherence and the presence of (quasi)off-diagonal long-range order, i.e. the existence of Bose-Einstein condensation according to the Penrose Onsager criterion [64]. Its Fourier transform yields the momentum distribution function, giving information on the velocity of the quantum particles in a given quantum state. The full space-time correlation function describes the evolution of the quantum system when removing or adding a particle to it. In space-time, it allows one to study the spread of correlations, and to test the Lieb-Robinson bound [65], and its generalisation to open quantum systems [66]. Its Fourier transform in frequency-wavevector space provides the spectral function, whose poles give the dispersion of single-particle excitations. The spectral function is also needed to describe the transport properties among two reservoirs across a quantum channel within linear response [67].
@@
In this work, we study the two-point correlation function of an interacting bosonic open quantum system described by the Bose-Hubbard model subjected to incoherent drive and losses. At non-zero interactions, similarly to its closed-system counterpart, this model is not integrable. We choose the case of spatially uniform pump and losses, where the non-equilibrium steady-state (NESS) density matrix is exactly known [68], and it is closely related to an infinite-temperature state at fixed chemical potential, independently of the interaction strength. Despite the steady state taking a simple form, we find that the two-point correlation function shows a non-trivial dependence on the interaction strength, displaying in particular typical temporal oscillations with frequency related to the interaction strength as well as an interaction-dependent exponential decay. We compare these results to the exactly known non-interacting limit as well as to the strongly interacting limit, where the dynamics of the different sites decouples and can be described using the exact results for the dissipative Kerr resonator [69]. We then proceed in obtaining the full spectral function, where in particular we find a doublon-like branch at excitation energy $U$ as well as a cosine-like excitation branch in the single-particle-hole sector. The features of the spectral function are a clear signature of the non-equilibrium nature of the quantum system, further supported by a perturbative calculation of the self-energy in the Keldysh formalism, and by the comparison with the closed-quantum system spectral function [70].
@@
The paper is organised as follows. We start by presenting the model in Sec. 2 and characterising the non-equilibrium steady state in Sec. 3. We report our analytical results for the two-point function obtained within the Keldysh formalism in Sec. 4. We then proceed to present our numerical results for the space-time correlations in Sec. 5, where we display the time-resolved correlations and analyse their decay. The spectral function is computed numerically and discussed in Sec. 6. Finally, Sec. 7 presents our concluding remarks.
@@ 2.1 Lindblad equation
$\hat{H}=\sum_{i=1}^{L}\Big{[}-J\big{(}b_{i}^{\dagger}b_{i+1}+\text{h.c.}\big{)}+\frac{U}{2}b_{i}^{\dagger}b_{i}^{\dagger}b_{i}b_{i}\Big{]}\;,$ (1)
@@
$\partial_{t}\rho=\mathcal{L}[\rho]=-\text{i}[\hat{H},\rho]+\gamma_{l}\sum_{i=1}^{L}\big{(}b_{i}\rho b_{i}^{\dagger}-\frac{1}{2}\{b_{i}^{\dagger}b_{i},\rho\}\big{)}+\gamma_{p}\sum_{i=1}^{L}\big{(}b_{i}^{\dagger}\rho b_{i}-\frac{1}{2}\{b_{i}b_{i}^{\dagger},\rho\}\big{)}\;.$ (2)
@@ 2.2 Correlation functions
where delimited-⟨$\langle$..$\rangle$ … ⟩ indicates the trace weighted with the respective density matrix, $\Theta(t)$ is the Heavyside step function, and $b_{j}(t)$ ($b^{\dagger}_{j}(t)$) is the bosonic annihilation (creation) field operator in the Heisenberg picture, time-evolved according to Lindbladian dynamics [^1] In the case of open quantum systems, a way of defining the time evolution of an operator $A$ is given by the adjoint superoperator $\bar{\mathcal{L}}$, according to $A(t)=e^{\bar{\mathcal{L}}t}[A]$ where ¯$\bar{\mathcal{L}}[A]=\text{i}[H,A]+\gamma_{l}\sum_{i=1}^{L}\big{(}b_{i}^{\dagger}Ab_{i}-\frac{1}{2}\{b_{i}^{\dagger}b_{i},A\}\big{)}+\gamma_{p}\sum_{i=1}^{L}\big{(}b_{i}Ab_{i}^{\dagger}-\frac{1}{2}\{b_{i}b_{i}^{\dagger},A\}\big{)}$. Notice that ${\rm Tr}\big{\{}A(t)\rho(0)\big{\}}={\rm Tr}\big{\{}A\rho(t)\big{\}}$. (in the open case), or Hamiltonian one (in the closed case).
@@
For a closed system the corresponding correlation function evaluated on the ground state $|\Psi_{0}\rangle$, which reads
$\Big{\langle}b_{j}^{\dagger}(t)b_{0}\Big{\rangle}_{\text{GS}}\equiv\langle\Psi_{0}|\text{e}^{\text{i}Ht}\,b_{j}^{\dagger}\text{e}^{-\text{i}Ht}\,b_{0}|\Psi_{0}\rangle\;,$ (6)
and evaluated on a finite temperature equilibrium ensemble, which takes the form
$\Big{\langle}b_{j}^{\dagger}(t)b_{0}\Big{\rangle}_{\text{equi}}\equiv\text{Tr}\Big{\{}\text{e}^{\text{i}Ht}\,b_{j}^{\dagger}\text{e}^{-\text{i}Ht}\,b_{0}\rho_{\text{equi}}\Big{\}}\;,$ (7)
where $\rho_{\text{equi}}$ is the canonical $\rho_{C}(\beta,N)=\text{e}^{-\beta\hat{H}}\,/{\rm Tr}[\text{e}^{-\beta\hat{H}}\,]$ or grand-canonical density matrix $\rho_{GC}(\beta,\mu)=\text{e}^{-\beta(\hat{H}-\mu\hat{N})}\,/{\rm Tr}[\text{e}^{-\beta(\hat{H}-\mu\hat{N})}\,]$ with $\beta^{-1}=k_{B}T$.
@@
for a homogeneous system.
@@ 3 Exact properties of the non-equilibrium steady state
3.1 Steady-state density matrix
@@
We remark that, in this idealised model, there is no high-energy cutoff on the pump intensity. We refer to Ref. [68] for a model with frequency dependent effective jump operators. The infinite dimensional Hilbert space of the model as written above renders the Hamiltonian unbounded. We point out that in the numerical calculations, we have to introduce a cutoff for the local occupation, which gives an upper bound for the Hamiltonian and ensures that the time evolution is a CPTP map as well as the existence of a NESS [52]. We have generalised the solution (9) for $\rho_{\text{NESS}}$ to the case of driven-dissipative free fermions and hard-core bosons on the lattice. The results are given in Appendix C and Appendix D respectively.